Determine the area corresponding to the following z-scores.

1. greater than 𝑧 = 2.31

2. less than 𝑧 = 1.20

3. between 𝑧 = βˆ’1.96 and 𝑧 = 2



1
Expert's answer
2021-03-26T07:57:27-0400

(1). P(z>2.31)= 1- 0.9896 = 0.0104






(2.) P(z<1.20)= 0.8849





(3.) P(-1.96<z<2)= 0.97725- 0.02500Β = 0.95225